# The Influence of Human-inspired Agentic Sophistication in LLM-driven Strategic Reasoners

### v1

@@ H S H S H S H S
$c_{\text{sim}},m_{0}$ 16.82 18.6 18.47 20.70 23.65 16.37 18.60 17.40
$c_{\text{sim}},m_{1}$ 18.74 16.98 18.16 21.04 14.68 15.84 17.33 18.02
$c_{\text{bio}},m_{0}$ 20.39 18.63 18.01 21.50 25.36 14.31* 18.82 18.28
$c_{\text{bio}},m_{1}$ 16.10 16.93 15.69* 23.14 22.69 15.59 17.22 17.54
In the case of Haiku, we observe that the inclusion of MoA has an overall negative correlation with the Wasserstein distance. Similar to the previous set of evaluations, the effects of the different configurations and the more complex reasoning model on Sonnet are inconsistent; no well-defined patterns emerge. The relative overoptimization of Sonnet and its poorer instruction following ability leads to a similar resistance to increasing complexity as we have seen in the experiments without profile specification. In 3/4 cases, Haiku and the $c_{\text{sim},m_{1}}$ configuration achieved the closest results. In the case where Haiku did worse, it produced an outlier distribution relative to its overall performance, which has a significantly lower variance and higher mean. The statistics show a high number of agents choosing 22 and 33, which are perfectly aligned with the $k=1$ and $k=2$ levels associated with guessing games. A manual random-sampled analysis of the produced reasoning samples indicates that the specific configuration seemingly caused the agent to pick up on the corresponding game-theoretic material. An in-depth analysis of the corresponding natural language reasoning samples may reveal further insights and is needed to validate our suspicions, but lies outside the scope of this paper.
We investigate how the zero guesses are spread over the subpopulations. The sparsity of zero guesses remains, but a consistent pattern we can observe is that, according to our expectation, expert agents guess zero more frequently than student agents, which mirrors human behaviour, where experts were thrice as likely to guess zero.
[…]
@@
Agent $-3.80^{***}$ $-1.69$ $0.25$ $-6.90^{***}$ $0.66$ $-12.00^{**}$
LLM $0.97^{***}$ $1.89$ $-0.25$ $2.46^{***}$ $-3.13^{*}$ $-5.50$
$C$ $-0.10$ $-0.11$ $0.25$ $0.73^{*}$ $-0.99$ $-2.00$
$M$ $-0.03$ $-0.79$ $0.25$ $1.51^{***}$ $-1.73$ $-5.50$
$\rho$ $0.19^{***}$ $0.01$ $-0.18$ $-0.08^{*}$ $0.03$ $0.21$
The group-specific regression analysis in Table 5 supports our earlier observations. Despite the sparsity of zero-guessing behaviour, we find that Reasoner agents ($R$) align significantly better with expert zero-guess rates. For distributional similarity (Wasserstein distance), most factors have no significant effect; the exception is the LLM model choice, which exhibits a significant effect for experts and a near-significant trend for students ($p=0.07$), with opposite signs across the two groups. Agent type and LLM model have a strong and consistent influence on $k$-level approximation for both cohorts: more complex reasoning agents using the smaller LLM (Haiku) achieve lower error. In contrast, contextual ($C$) and instruction complexity ($M$) have only minor and non-significant effects for students, while for experts, increased complexity correlates with higher $k$-level error. Finally, the Spearman $\rho$ correlation analysis reveals a significant but modest association between overall sophistication and $k$-level error: positive for students, suggesting over-complexity harms alignment, and negative for experts, indicating modest improvements. No monotonic trends emerge for the other two metrics.
@@ 4.3 Out-of-Sample Validation
Most guessing game experiments involve $n$ players selecting numbers between 00 and $100$ [29]. In contrast, our 2-player variant may be underrepresented or entirely absent from LLM training. We treat our setup as an out-of-sample case and further examine generalization by shifting key parameters to diverge from training data priors.
@@
Assuming human understanding of the game is invariant to the number range, we normalise shifted-range guesses by subtracting 100 and compare them to standard-range distributions. We then compute Wasserstein distances and t-statistics between original and scaled conditions to identify robustness across agent variations.
@@
$c_{\varnothing},m_{0}$ 2.70 0.02 15.99 12.34* 7.69 2.77* 13.20 18.36*
$c_{\varnothing},m_{1}$ 9.06 2.10* 9.26 6.39* 13.62 5.87* 27.84 -0.55
$c_{\text{sim}},m_{0}$ 7.21 4.70* 12.19 8.61* 7.60 1.54 13.63 14.41*
$c_{\text{sim}},m_{1}$ 8.81 -4.06* 12.64 8.33* 14.76 7.00* 14.78 5.12*
$c_{\text{bio}},m_{0}$ 5.48 5.46* 14.05 12.08* 8.39 1.04 12.29 15.03*
$c_{\text{bio}},m_{1}$ 7.34 4.38* 13.79 11.80* 13.38 6.31* 12.36 26.41*
The results in Table 6 reinforce earlier findings: Haiku does better than Sonnet, and $R$ agents are more dissimilar. While increasing the $C$ complexity does not lead to significant and consistent changes, $m_{1}$ increases the distance score in each case over $m_{0}$.
@@ 4.4 Cost of Complexity
$c_{\varnothing},m_{0}$ $c_{\varnothing},m_{1}$ $c_{\text{sim}},m_{0}$ $c_{\text{sim}},m_{1}$ $c_{\text{bio}},m_{0}$ $c_{\text{bio}},m_{1}$
@@ 5 Conclusions
The first author would like to express his gratitude to David Levine for his valuable insights.

### v2

@@ H S H S H S H S
$c_{\text{sim}},m_{0}$
16.82 18.6 18.47 20.70 23.65 16.37 18.60 17.40
$c_{\text{sim}},m_{1}$
18.74 16.98 18.16 21.04 14.68 15.84 17.33 18.02
$c_{\text{bio}},m_{0}$
20.39 18.63 18.01 21.50 25.36 14.31* 18.82 18.28
$c_{\text{bio}},m_{1}$
16.10 16.93 15.69* 23.14 22.69 15.59 17.22 17.54
Equilibrium Play
We next investigate how the zero guesses are spread over the subpopulations. As Table 5, although overall rates remain sparse, experts predictably guess zero more frequently than students, consistent with the human data, where experts were three times more likely to play the dominant strategy.
Table 5: Zero guesses per specified profile. Bold values are the cohort-wise best match ($11.9$ and $37.69$) per agent type, italic values denote the best match per LLM and * marks the best score in the table.
Config Student Expert
Simple Reasoner Simple Reasoner
H S H S H S H S
$c_{\text{sim}},m_{0}$
0 0 0 0 0 10 0 0
$c_{\text{sim}},m_{1}$
0 0 0 0 28 20 0 0
$c_{\text{bio}},m_{0}$
0 0 0 2 0 18 0 0
$c_{\text{bio}},m_{1}$
0 0 0 0 0 22 0 2
Factor Analysis
Table 6: Cohort-specific regression coefficients and Spearman correlations for sophistication influence across three behavioural alignment metrics. Stars indicate statistical significance: * $p<.05$, ** $p<.01$, *** $p<.001$.
@@
Agent $3.80^{***}$ 1.69-1.69 $0.25$ $6.90^{***}$ $0.66$ $12.00^{**}$
LLM $0.97^{***}$ $1.89$ 0.25-0.25 $2.46^{***}$ $3.13^{*}$ 5.50-5.50
$C$ 0.10-0.10 0.11-0.11 $0.25$ $0.73^{*}$ 0.99-0.99 2.00-2.00
$M$ 0.03-0.03 0.79-0.79 $0.25$ $1.51^{***}$ 1.73-1.73 5.50-5.50
$\rho$ $19^{***} 0.0$10.01 0.18-0.18 $08^{*} 0.0$30.03 $0.21$
Table 6 summarises regression coefficients and Spearman correlations for sophistication factors across three behavioural metrics. The results support earlier observations. Despite the sparsity of zero-guessing behaviour, we find that $R$ agents align significantly better with expert zero-guess rates. For distributional similarity (Wasserstein distance), most factors show no significant effect, with the notable exception of LLM choice, which strongly affects experts and shows a trend for students ($p=0.07$). Agent type and LLM model have a strong and consistent influence on $k$-level approximation for both cohorts: more complex reasoning agents using the smaller LLM (Haiku) achieve lower error. In contrast, contextual ($C$) and instruction complexity ($M$) have only minor and non-significant effects for students, while for experts, increased complexity correlates with higher $k$-level error. Finally, the Spearman $\rho$ correlation analysis reveals a significant but modest correlation between overall sophistication and $k$-level error: positive for students, suggesting over-complexity harms alignment, and negative for experts, indicating modest improvements. No monotonic trends emerge for distributional similarity or equilibrium play.
@@ 4.3 Out-of-Sample Validation
Most traditional guessing game experiments involve $n$ players selecting numbers between 0 and $100$ [29]. In contrast, our 2-player variant may be underrepresented or entirely absent from LLM training. We treat our setup as an out-of-sample case and further examine generalization by shifting key parameters to diverge from training data priors.
@@
Assuming human understanding of the game is invariant to the number range, we normalise shifted-range guesses by subtracting 100 and compare them to standard-range distributions. We then compute Wasserstein distances and t-statistics between original and scaled conditions to identify robustness across agent types and configurations. The results in Table 7 reinforce earlier findings: Haiku does better than Sonnet, and $R$ agents are more dissimilar. While increasing the $C$ complexity does not lead to significant and consistent changes, $m_{1}$ increases the distance score in each case over $m_{0}$.
@@
$c_{\varnothing},m_{0}$
2.70 0.02 15.99 12.34* 7.69 2.77* 13.20 18.36*
$c_{\varnothing},m_{1}$
9.06 2.10* 9.26 6.39* 13.62 5.87* 27.84 -0.55
$c_{\text{sim}},m_{0}$
7.21 4.70* 12.19 8.61* 7.60 1.54 13.63 14.41*
$c_{\text{sim}},m_{1}$
8.81 -4.06* 12.64 8.33* 14.76 7.00* 14.78 5.12*
$c_{\text{bio}},m_{0}$
5.48 5.46* 14.05 12.08* 8.39 1.04 12.29 15.03*
$c_{\text{bio}},m_{1}$
7.34 4.38* 13.79 11.80* 13.38 6.31* 12.36 26.41*
@@ 4.4 Cost of Complexity
$c_{\varnothing},m_{0}$
$c_{\varnothing},m_{1}$
$c_{\text{sim}},m_{0}$
$c_{\text{sim}},m_{1}$
$c_{\text{bio}},m_{0}$
$c_{\text{bio}},m_{1}$
@@ 5 Conclusions
The authors would like to express their gratidude to David Levine for his constructive comments and fruitful discussions about the game-theoretic material and the EWA model.